# XpertAI: uncovering model strategies for sub-manifolds

### v1

@@ Abstract
In recent years, Explainable AI (XAI) methods have facilitated profound validation and knowledge extraction from ML models. While extensively studied for classification, few XAI solutions have addressed the challenges specific to regression models. In regression, explanations need to be precisely formulated to address specific user queries (e.g. distinguishing between ‘why is the output above 0?’ and ‘why is the output above 50?’). They should furthermore reflect the model’s behavior on the relevant data sub-manifold. In this paper, we introduce XpertAI, a framework that disentangles the prediction strategy into multiple range-specific sub-strategies and allows formulating precise queries about the model (the ‘explanandum’) as a linear combination of those sub-strategies. XpertAI is formulated generally to work alongside popular XAI attribution techniques, based on occlusion, gradient integration, or reverse propagation. Qualitative and quantitative results, demonstrate the benefits of our approach.
Machine Learning, ICML
Figure 1: Left: Methodological overview of our proposed XpertAI approach. We introduce a virtual layer of range experts to get insights into the model behavior on output-range specific sub-manifolds. In the forward pass, we retain the model behavior through the additive structure of experts (top, right). At the same time, they represent a well-organized latent space that can be systematically queried when explaining the model (bottom left). Right: While the naive application of attribution methods typically answers questions from a generic point of view (grey) our approach enables answers to more nuanced queries as defined by the user (red). For the task of wine quality prediction, we see significant structural changes in the explanation for different reference values (center). When analyzing the losses of a wind turbine, contextualized attributions correspond more faithfully to the true losses (right).
@@ 1 Introduction
Machine learning has provided powerful predictive models for numerous scientific and industrial applications. As the use of ML models for critical autonomous decisions increases, there is a growing demand for establishing trust while maintaining their predictive capabilities. Explainable artificial intelligence (XAI) has emerged as a step towards enhancing transparency and allows for insights into the inner workings of these highly complex AI models (Arrieta et al., 2020; Samek et al., 2021). XAI can be utilized for both, model validation against expert intuition and for obtaining new insights into the data-generating processes under investigation (Krenn et al., 2022; Klauschen et al., 2024).
@@
So far, the predominant focus within XAI has been placed on understanding the decisions made by classification models (Baehrens et al., 2010; Zeiler & Fergus, 2014; Bach et al., 2015; Zhou et al., 2016; Ribeiro et al., 2016). The widely used family of post-hoc attribution methods aims to achieve this by allocating evidence for a particular class across the corresponding input features. Thus they indicate the extent to which each feature has contributed to the model output. In this process, the model’s decision boundary serves as a natural point of reference for the explanation.
In regression, on the other hand, the matter is more complex. There, the equivalent to the decision boundary needs to be defined for every single query, since it is a priori unknown which of the two questions ‘why is the output above 50?’ or ‘why is the output above 0?’ is most relevant for the user (Letzgus et al., 2022). Moreover, in non-linear problems, sub-manifolds on which the model builds specific responses are to be expected, for example, for different output values: Consider, for example, the task of predicting the quality of wine based on some descriptive features (year, grape, price, etc.). Intuitively, the explanation of what makes a premium wine better than a decent alternative is likely to differ significantly from the explanation of what makes it better than a low-quality rotgut, which should be reflected in the attributions.
@@
To address these challenges, we propose our XpertAI framework. The basic idea is to decompose the regression model output into a set of additive basis functions, so-called range experts (compare Figure 1). Each range expert is dedicated to capturing the model behavior within a specific, output-range-dependent sub-manifold. Subsequently, range experts can be queried by the user with any state-of-the-art attribution method to obtain explanations that are contextualized to the individual explanatory needs. We demonstrate the benefits of our method on several (controlled and real-world) problems (see Figure 1). Regarding the wine example from above, we indeed find that the model considered the price the most important input feature to distinguish an excellent wine from a bad one. But when explaining with respect to decent alternatives (close-by-reference values), other quality-related features become much more important. In another case study, we use attributions to monitor the performance of a wind turbine. There, we find that our contextualized explanations more faithfully capture the performance losses, which enables better maintenance decisions in practice. In addition to these qualitative insights, we report improved faithfulness through better contextualization with XpertAI.
@@ 2 Related Work
Our proposed method relates to several specific areas of XAI, which we will briefly discuss below (see e.g. (Samek et al., 2021; Arrieta et al., 2020) for XAI reviews).
@@ 2.1 Mixture of Experts
The concept of Mixture of experts (MoE) has been used for several decades for the purpose of learning better models (Jacobs et al., 1991; Hampshire & Waibel, 1992; Pawelzik et al., 1996). Beyond model improvement, recent work has demonstrated the value of the MoE approach for achieving more transparent models by combining inherently interpretable linear experts (Ismail et al., 2023). In contrast, our approach utilizes MoEs for explaining models in a post-hoc manner, without restriction on the structure of the model, and steering the expert to become ‘range experts’ focusing on specific value ranges.
@@ 2.2 Context in XAI attribution methods
Generally speaking, every explanation requires context to be meaningful. This also applies to XAI attribution methods which incorporate a baseline in input space $\widetilde{x}$ or latent space $\widetilde{a}$ to do so. Occlusion-based methods, such as Shapley values (Lundberg & Lee, 2017), for example, need the user to select a baseline in input space to occlude with. Gradient Integration methods (Sundararajan et al., 2017) need to define a start of the integration path and deep Taylor expansions (Montavon et al., 2017) root points. Propagation-based methods allow for a baseline selection in latent-space, either explicitly (Shrikumar et al., 2017) or implicitly via the choice of propagation rules (Montavon et al., 2017; Bach et al., 2015). If we have a concrete counterfactual sample $x^{\prime}$ it can therefore be utilized to obtain a contextualized result (Albini et al., 2022; Dombrowski et al., 2023). In practice, however, this is often not the case which leaves the selection to the user.
Finding an appropriate baseline or root point is a difficult task. At the same time, it has considerable influence on the explanation (Sundararajan & Najmi, 2020). Suitable candidates have to be on the data manifold, preferably in proximity to the sample in question (Dombrowski et al., 2023). In classification, a baseline additionally has to lay on the decision boundary to obtain meaningful attributions. In regression, each $\widetilde{x}$ inherently corresponds to a reference value $f(\widetilde{x})=\widetilde{y}$ on the model output scale, which adds yet another degree of freedom to be considered.
One solution that comes to mind is to compute expected attributions across multiple baselines (Kapishnikov et al., 2019; Lundberg & Lee, 2017). For example, by conditionally drawing from the dataset. The conditional mean around some desired reference value ($x^{\prime}=\mathbb{E}(x|\widetilde{y})$) could also serve as a synthetic counterfactual baseline, although $f(x^{\prime})$ usually does not coincide with the exact desired $\widetilde{y}$ for non-linear models. Alternatively, Bayesian- or generative modelling offer a computationally more expensive way to synthesize a counterfactual (Van Looveren et al., 2021; Spooner et al., 2021). In this contribution, we propose a practical solution that ensures contextualization by design and therefore increases robustness against suboptimal baselines in input space.
@@ 2.3 Disentangled XAI and Virtual Layers
While refining the question to be asked is essential in a regression setting (see previous sections), many works have focused on independently refining the explanation itself (mainly in a classification context). Specifically, enriching explanations by identifying its multiple components, associated with distinct abstract concepts. These can be obtained in a supervised manner (Kim et al., 2018; Zhao et al., 2023), in an unsupervised manner (Vielhaben et al., 2023; Chormai et al., 2022), or by directly inspecting neurons (Zeiler & Fergus, 2014; Zhou et al., 2016; Achtibat et al., 2023).
This kind of analysis often involves an informed transformation of latent representations to obtain a meaningful or relevant ’concept space’, followed by the inverse transformation to leave the overall model behaviour intact. Therefore, these approaches are referred to as virtual layers. (Chormai et al., 2022), for example, extract sub-concepts that jointly contribute to the explanation of an overall class concept. Likewise, (Vielhaben et al., 2024) generates a Fourier basis on which the prediction of speech samples can be more efficiently analyzed, and (Linhardt et al., 2024) introduces a virtual PCA layer, which disentangles verified from unverified factors of variation and subsequently prune the latter for increased robustness.
We extend these prior efforts to the broad domain of regression, by introducing a novel technique aiming at disentangling global phenomena that exert influence consistently across the entire range of potential regression outputs from more localized, context-specific patterns (see Figure 1).
@@
3 Our Method: XpertAI
In the following, we introduce our novel method, called XpertAI for explaining neural network regression models. Our approach is inspired by the MoE concept and consists of appending range experts to a given ML model, thus allowing the user to formulate precise queries for which range they need an explanation. This appendage can be seen as a virtual layer inserted in the neural network, which – while leaving the overall prediction function intact – enriches it by providing the basis for query formulation and explanation.
@@ 3.1 Adding Range Experts
where $\rho_{0,\tau}(y)=\min(\max(y,0),\tau)$ clips the input to the interval $[0,\tau]$. A low $\tau$ corresponds to a more specialized expert. The transformation in Eq. (5) is also known as thermometer coding. The architecture that results from appending these experts is shown in Fig. 2. Assuming the values of $y$ are always positive (which we can ensure through offsetting) we can reconstitute the output prediction by summing the experts’ outputs:
@@
The mapping from $y$ to $z$ and back to $y$ can be seen as a virtual layer whose insertion does not affect the input-output mapping, but that provides additional functionality. Unlike previous formulations of virtual layers (Vielhaben et al., 2024), ours is placed at the output, enabling a disentanglement of the explanation in terms of those range experts.
@@ 3.2 Querying / Explaining
Whereas the disentanglement performed above provides a more detailed view of the prediction behavior than a simple explanation, the user is often interested in particular aspects of it. Our approach lets the user formulate a query (or ‘explanandum’) as a linear combination of the range experts:
@@
An example of such a query is given in Fig. 2 (right). For example, if the user is interested in what makes a prediction $y=60$ larger than a reference value of $50$, the query $q$ can be shaped in the form of a sigmoid centered at the reference value $50$.
@@
Note that for explanation techniques that fulfill the linearity axiom w.r.t. the last layer of representation, we can further develop the expression of the explanation as:
@@
These two results are easily retrievable by observing the specific structure of the explanation given in Eq. (9). Proofs can be found in Appendix A.
@@ 3.3 Structural Disentanglement
When the underlying explanation method relies not directly on the ML model’s output but on its computational graph (e.g. LRP), the latter must be disentangled. Clearly, such a structural disentanglement is missing as the mapping from activations $a$ to the expert’s outputs $z$ passes through a one-dimensional bottleneck $y$ (the original real-valued output). We propose to replace the original mapping $a\mapsto(z_{m})_{m}$ by a surrogate model
@@
which encourages that the surrogate’s output is correct within-range and on the correct side outside-range. We then solve $\min_{\theta}\mathbb{E}[\sum_{m}\ell(s_{m},z_{m})]$ with $\mathbb{E}[\cdot]$ denoting the expectation over the training data. To preserve not only the prediction output of the original model but also its prediction strategy (i.e. the feature it uses) further steps are needed. One approach is to enforce the loss function not only on the data but also on perturbations of the data (Stutz et al., 2019). For example, activations can be randomly turned off (with a probability chosen between 00 and $1$). This perturbation scheme ensures in particular that the Shapley value explanations of the original and disentangled models become similar (i.e. that they predict the same for the same reasons). Furthermore, we find that freezing the bias in the output layer is important in order to achieve the desired structural disentanglement.
4 XpertAI-Opinion: insights into model behavior on sub-manifolds
We now demonstrate how our XpertAI approach can help users disentangle local and global effects for meaningful insights in three different case studies. For each of them, we briefly introduce the dataset, model and XpertAI setting, before presenting the insights. We present results from using both, Integrated Gradients and LRP. Details on all case studies can be found in Appendix C.
@@ 4.1 Uncovering output-scale-specific strategies
We start this investigation in a rather controlled environment. For that, we adopt the world’s arguably best-known classification data set, MNIST (Deng, 2012), and transform it into a regression problem (rMNIST). For simplicity, we take the subset of only three digits (4,5, and 6) and calculate a one-dimensional t-SNE representation (van der Maaten & Hinton, 2008), which henceforth serves as a new label for each sample. Additionally, we ensure labels are distributed uniformly between values of zero and one. As a result, the individual digits populate continuous parts of the output dimension (in our case sorted by digit magnitude, which facilitates interpretation) while sorting within each digit bin is based on the respective digit’s rotation (compare Figure 3). We now train a vanilla CNN model architecture to learn this mapping from image to output scale. For contextualized insights, we train three range experts (one for each digit range). We first discuss qualitative results and present its quantitative evaluation in Section 5.
Figure 4 shows a comparison between the standard and the XpertAI explanations for both, IG and LRP. In the bottom row, we show the average naive attributions over all samples within the respective output range. The explanations from the lowest-range expert (second row from below) are very similar to the original ones since they implicitly explain relative to the same reference value. The range experts of the upper digit ranges (marked in colors), however, enable more granular insights. It is visible how the range experts for the output range populated with the digits five and six (marked in orange and blue) seem to focus specifically on the rotation of the digit. A rotation to the right is associated with lower values (negative attribution, blue) and vice versa. See Appendix C for more basis functions. In Section 5, we will see that these qualitative differences in attributions also result in improved quantitative evaluation scores for attribution faithfulness.
@@
Figure 4: Mean attributions over different cohorts of samples (columns) and basis functions (rows). The bottom row represents naive attributions. The top three rows correspond to ascending range-specific XpertAI bases. Note, how only the latter exposes the digit rotation within the respective digit ranges (orange/blue).

### v3

# XpertAI: uncovering regression model strategies for sub-manifolds

@@ Abstract
In recent years, Explainable AI (XAI) methods have facilitated profound validation and knowledge extraction from ML models. While extensively studied for classification, few XAI solutions have addressed the challenges specific to regression models. In regression, explanations need to be precisely formulated to address specific user queries (e.g. distinguishing between ‘why is the output above 0?’ and ‘why is the output above 50?’). They should furthermore reflect the model’s behaviour on the relevant data sub-manifold. In this paper, we introduce XpertAI, a framework that disentangles the prediction strategy into multiple output range-specific sub-strategies and allows the formulation of precise queries about the model as a linear combination of those sub-strategies. XpertAI is formulated generally to work alongside popular XAI attribution techniques, based on occlusion, gradient integration, or reverse propagation. Qualitative and quantitative results demonstrate the benefits of our approach.
Keywords:
XAI Post-hoc attributions Regression Mixture of experts Contrastive explanations
@@ 1 Introduction
Machine learning has provided powerful predictive models for numerous scientific and industrial applications. As the use of ML models for critical autonomous decisions increases, there is a growing demand for establishing trust while maintaining their predictive capabilities. Explainable artificial intelligence (XAI) has emerged as a step towards enhancing transparency and allows for insights into the inner workings of these highly complex AI models [5, 40]. XAI can be utilized for both, model validation against expert intuition as well as for obtaining new insights into the data-generating processes under investigation [24, 23].
Figure 1: Top: Conceptual overview of our proposed XpertAI approach. We add a layer of range expert neurons, each responsible for mimicking the original model behaviour on a range-specific sub-manifold of the data. During inference, the outputs of all range experts are added up and result in the original model output. When explaining, we isolate output-range-specific effects by querying only the respective or a combination of range experts. Bottom: While the naive application of attribution methods typically answers questions from a generic point of view (grey) our approach enables answers to more nuanced questions as defined by the user (red). For the tasks of wine quality prediction and attributing losses of a wind turbine, we see significant structural changes in the explanations. For details see sections 4.2 and 4.3.
@@
So far, the predominant focus within XAI has been placed on understanding the decisions made by classification models [7, 49, 6, 51, 38]. The widely used family of post-hoc attribution methods aims to achieve this by allocating evidence for a particular class across the corresponding input features. In doing so, they indicate the extent to which each feature has contributed to the model output. In this process, the model’s decision boundary serves as a natural point of reference for the explanation. In regression, on the other hand, the equivalent to the decision boundary needs to be defined for every single query, since it is a priori unknown which of the two questions ‘why is the output above 50?’ or ‘why is the output above 0?’ is most relevant for the user [26]. Moreover, in non-linear problems, sub-manifolds on which the model builds specific responses are to be expected, for example, for different output values.
@@
To address these challenges, we propose our XpertAI framework. The basic idea is to decompose the output of the regression model into a set of additive basis functions, the so-called range experts (compare Figure 1, top). Each range expert is dedicated to capturing the model behavior within a specific, output-range-dependent sub-manifold. Subsequently, the user can query the range experts with any state-of-the-art attribution method to obtain explanations that are contextualized to the individual explanatory needs. We demonstrate the benefits of our method on several (controlled and real-world) problems (see Figure 1, bottom). We, for example, find that a model considered the price the most important input feature to distinguish an excellent wine from a bad one. But when explaining with respect to decent alternatives (close-by-reference values), other quality-related features become much more important. In another case study, we used attributions to monitor the performance of a wind turbine. There, we find that our contextualized explanations more faithfully capture the performance losses, which enables better maintenance decisions in practice. In addition to these qualitative insights, we report improved faithfulness through better contextualization with XpertAI. An implementation is available online. [^1] https://github.com/sltzgs/XpertAI
@@ 2 Related Work
Our proposed method relates to several specific areas of XAI, which we will briefly discuss within this chapter (see e.g. [40, 5] for XAI reviews).
@@ 2.1 Mixture of Experts
The Mixture of Experts (MoE) framework [20, 14, 35] follows a divide-and-conquer strategy, commonly used to enhance model performance. Recent work has applied MoEs for transparency by combining interpretable linear experts [19]. In contrast, our approach utilizes MoEs for explaining models in a post-hoc manner, without restriction on the structure of the model, and steering the expert to become ‘range experts’ focusing on specific value ranges. This is achieved by dividing the data into sub-manifolds according to the output range of a regression model, a way of domain-informed gating, and explaining the model strategy within these specific regions.
@@ 2.2 Context in XAI attribution methods
Generally speaking, every explanation requires context to be meaningful. When explaining the outcome of a classification model, the decision boundary serves as a natural point of reference. Contrastive explanations have been proposed to better incorporate user-specific context into the explanation [21, 43, 28]. For regression models, on the other hand, explanations depend on the reference output relative to which we seek an explanation [26]. XAI attribution methods allow for the incorporation of context through baselines, which depending on the method have to be chosen in input space [29, 45] or latent space [41, 31, 26]. Each baseline then corresponds to a respective reference value ($\widetilde{y}$). In practice, the choice of baselines represents a challenge with fundamental impact on the outcome of the explanation. In this work, we therefore propose a practical solution that ensures contextualization by design for regression models and, as a result, increases robustness against suboptimal baseline choices.
@@ 2.3 Disentangled XAI and Virtual Layers
While refining the question to be asked is essential in a regression setting, many works have focused on independently refining the explanation itself (mainly in a classification context). Specifically, enriching explanations by identifying its multiple components, associated with distinct abstract concepts. These can be obtained in a supervised manner [22, 50], in an unsupervised manner [46, 9], or by directly inspecting neurons [49, 51, 2]. This kind of analysis often involves an informed transformation of latent representations to obtain a meaningful or relevant ’concept space’, followed by the inverse transformation to leave the overall model behaviour intact [48]. Therefore, these approaches are referred to as virtual layers. [9], for example, extract sub-concepts that jointly contribute to the explanation of an overall class concept. Likewise, [47] generates a Fourier basis on which the prediction of speech samples can be analyzed more efficiently, and [27] introduces a virtual PCA layer, which disentangles verified from unverified factors of variation and subsequently prune the latter for increased robustness. We extend these efforts to the broad domain of regression, by introducing a novel technique that aims to disentangle global phenomena that exert influence consistently across the entire range of potential regression outputs from more localized context-specific patterns (see Figure 1, top).
3 Our Method: XpertAI
In the following, we introduce our novel method, called XpertAI for explaining neural network regression models. Our approach is inspired by the MoE concept and consists of appending range experts to a given ML model, thus allowing the user to formulate precise queries for which range they need an explanation. This appendage can be seen as a virtual layer inserted in the neural network, which – while leaving the overall prediction function intact – enriches it by providing the basis for query formulation and explanation. Figure 2 conceptually depicts the method and its notation, with details in the following sections.
@@ 3.1 Adding Range Experts
where $\rho_{0,\tau}(y)=\min(\max(y,0),\tau)$ clips the input to the interval $[0,\tau]$. A low $\tau$ corresponds to more specialized experts. The kind of transformation in Eq. (5) is also known as thermometer coding. The architecture that results from appending these experts is shown in Fig. 2. Assuming the values of $y$ are always positive (which we can ensure through offsetting) we can reconstitute the output prediction by summing the experts’ outputs:
@@
The mapping from $y$ to $z$ and back to $y$ can be seen as a virtual layer which does not affect the input-output mapping, but that provides additional functionality. Unlike previous formulations of virtual layers [47], ours is placed at the output, enabling a disentanglement of the explanation in terms of output ranges.
@@ 3.2 Querying / Explaining
Whereas the disentanglement performed above provides a more detailed view of the prediction behaviour than a simple explanation, the user is often interested in particular aspects of it. Our approach lets the user formulate a query (or ‘explanandum’) as a linear combination of the range experts:
@@
An example of such a query is given in Fig. 2 (right). For example, if the user is interested in what makes a prediction $y=60$ larger than a reference value of $50$, the query $q$ can be shaped in the form of a sigmoid centred at the reference value $50$.
@@
Note that for explanation techniques that fulfil the linearity axiom w.r.t. the last layer of representation, we can further develop the expression of the explanation as:
@@
These two results are easily retrievable by observing the specific structure of the explanation given in Eq. (9). Proofs can be found in Appendix 0.A.
@@ 3.3 Structural Disentanglement
When the underlying explanation method relies not directly on the ML model’s output but on its computational graph (e.g. LRP), the latter must be disentangled. Clearly, such a structural disentanglement is missing as the mapping from activations $a$ to the expert’s outputs $z$ passes through a one-dimensional bottleneck $y$ (the original real-valued output). We propose to replace the original mapping $a\mapsto(z_{m})_{m}$ by a learned surrogate model $(s_{m})$:
@@
which encourages that the surrogate’s output is correct within-range and on the correct side outside-range. We then solve $\min_{\theta}\mathbb{E}[\sum_{m}\ell(s_{m},z_{m})]$ with $\mathbb{E}[\cdot]$ denoting the expectation over the training data. To preserve not only the prediction output of the original model but also its prediction strategy (i.e. the feature it uses) further steps are needed. One approach is to enforce the loss function not only on the data but also on perturbations of the data [44]. For example, activations can be randomly turned off (with a probability chosen between 00 and $1$). This perturbation scheme ensures in particular that the Shapley value explanations of the original and disentangled models become similar (i.e. that they predict the same for the same reasons). Furthermore, we find that freezing the bias in the output layer is important to achieve the desired structural disentanglement.
3.4 XpertAI evaluation
We evaluate our proposed approach qualitatively (Section 4) and quantitatively (Section 5). In both cases we rely on either a (constructed) problem that allows for validation against some sort of ground truth, or the observation of model behaviour under attribution-guided, meaningful input perturbations. [13] proposed a regression-specific metric called the area between the curves (ABC). The ABC is defined as the area between the model output when occluding a sample’s features in the order of attribution magnitudes and a straight line connecting $f(x)$ and $f(x^{\prime})$ (which corresponds to random sorting). Since sorting ascending and descending can result in asymmetrical curves, we sum over both areas [8]. For a balanced result, we normalize by the distance between the sample and the baseline when averaging. Higher values of ABC are better.
Furthermore, the challenge of including context in attribution methods (Sec. 2.2) naturally extends to occlusion-based evaluation (ergo, what to occlude with?). To ensure that we evaluate attributions within the relevant output range of function $f(x)$, where we account for context-specific (local) effects, we occlude with a domain-specific counterfactual [3, 13]. Therefore, we sample conditional $x^{\prime}=D(x|y=\widetilde{y})$ from the available data set $D$ with which we then occlude and average the respective ABCs over multiple draws.
4 XpertAI-Opinion: insights into model behaviour on sub-manifolds
We now demonstrate how our XpertAI approach can help users disentangle local and global effects for meaningful insights in different case studies. First, we uncover output-scale-specific strategies for image regression problems (4.1). Then, we explain the quality of red wine (4.2) and the production losses of a wind turbine due to a technical malfunction (4.3). For each of the problems, we briefly introduce the dataset, model and XpertAI setting, before presenting the insights. We present results from using both, Integrated Gradients and Layer-wise Relevance Propagation (LRP). Details on all case studies can be found in Appendix 0.C.
@@ 4.1 Uncovering output-scale-specific strategies
First, we adopt the well-known MNIST [11] dataset and transform it into a regression problem (rMNIST). For simplicity, we take the subset of only three digits (4,5, and 6) and calculate a one-dimensional t-SNE representation [30], which henceforth serves as a new label for each sample. Additionally, we ensure labels are distributed uniformly between values of zero and one. As a result, the individual digits populate continuous parts of the output dimension (in our case sorted by digit magnitude, which facilitates interpretation) while sorting within each digit bin is based on the respective digit’s rotation (compare Figure 3). We now train a vanilla CNN model architecture to learn this mapping from image to output scale. For contextualized insights, we train three range experts (one for each digit range). We first discuss qualitative results and present its quantitative evaluation in Section 5.
@@
Figure 4: Mean attributions over different cohorts of samples (columns) and basis functions (rows). The bottom row represents naive attributions. The top row corresponds to the respective range-specific expert XpertAI bases. Note, how only the latter exposes the digit rotation within the digit ranges (orange/blue).
Figure 4 shows a comparison between the standard and the XpertAI explanations for both, LRP and IG. We contrast the average naive attributions over all samples within the respective output range (bottom row) with the explanations obtained with the respective range experts (top row). The explanations for the digit range 4 remain the same since both implicitly assume the same reference value (zero on the output scale). The expert attributions for the upper digit ranges (marked in orange and blue), enable more granular insights. It is visible how the range experts focus specifically on the rotation of the digit: a rotation to the right is associated with lower values (negative attribution, blue) and vice versa. See Appendix 0.C for more basis functions. In Section 5, we will see that these qualitative differences in attributions also result in improved quantitative evaluation scores for attribution faithfulness.
Now, let’s consider an illustrative regression task closer to real-world applications: biological age estimation from facial images [15, 4, 1] (see Appendix 0.C for model and data set details). Intuitively, the explanation for a person with a high age should be structurally different when being contrasted with a much younger age or an only slightly younger one. We, therefore, focus on a high-age cohort (individuals predicted to be above 77 years) and train three range experts ($\tau$ = 38.5 years). Figure 5, left, shows LRP attributions for the original model, averaged over the respective samples. Our proposed approach now allows us to disentangle these further using the respective age-specific basis functions. As expected, the explanation relative to the ’young’ basis (centre) overall contains much more positive evidence than the one with respect to the closer reference value of 77 years (right). Additionally, we can see that the latter is more fine-grained with a remaining focus on the person’s eyes and, surprisingly, we discovered a sign-flip for the oronasal region with a particular focus on lips and teeth. While the mouth has been reported to be an area particularly vulnerable to biases in age estimation from facial images [12], we put the faithfulness of these particular explanations to the test.
Figure 5: Comparison of average attributions for standard LRP (left) and two different XpertAI basis functions. Red indicates positive, and blue negative evidence. We can see that the disentangled explanations allow for much more fine-grained conclusions. Interestingly, the sign flip of the mouth area was masked by the strong attributions with respect to the original basis. We test for its faithfulness in Figure 6.
In Figure 6, we compare the effect of occluding the respective parts (eyes and mouth) of people’s faces with a generic average over all images. One example of each is shown at the right of the figure. Recall that this means we mask the eyes and mouth section with a relatively ’younger’ version. The chart shows the respective change in the model’s output. In line with intuition, and the explanations, age is indeed consistently decreased when occluding the eyes. Masking the mouth area with relatively ’younger’ mouths, however, indeed results in an increase of the model’s average prediction in many cases. The XpertAI-basis therefore constitutes the more faithful explanation since the attributions correctly captured the sign flip in model behaviour.
Figure 6: Validation of findings from disentangling age prediction (compare Fig. 5). We occlude relevant parts of the image according to the disentangled explanations (eyes and mouth) with the dataset-wide average face (two examples on the right). We then observe the effect on the model output relative to the original model prediction. For the high-age cohort, the two areas have distinctly different effects. Occluding the eyes with a relatively younger pair results in a consistent decrease in the predicted age. Occluding the mouth region, however, results in an increase for many of the samples. This model behaviour is in line with our insights from the disentangled explanations.
In conclusion, the disentangled basis explanations enabled more detailed insights into the model’s inner workings for both, the rMNIST and the age-prediction cases. They revealed effects that were not apparent from the naive explanations of the original model, since their highly aggregated nature did not allow for more fine-grained insights.